# MoTE: Mixture of Ternary Experts for Memory-efficient Large Multimodal Models

### v1

@@ 1 Introduction
Large Multimodal Models (LMMs) AAA+ (24); MGF+ (24); ZGG+ (24); WBT+ (24); CWT+ (24); BCL+ (25) have achieved remarkable performance across a wide range of downstream tasks, including visual question answering and autonomous computer agents. However, as model size increases, the rising inference cost presents significant challenges for deploying LMMs efficiently. To address this, Mixture-of-Experts (MoE) LLX+ (21); FZS (22); DLF+ (24) introduces a mechanism that maintains a large pool of experts while activating only a subset for each input, thereby improving computational efficiency. Although MoE models significantly reduce FLOPs, they generally have a higher memory footprint, making deployment on edge devices challenging. For example, when training multimodal MoE up-cycled from Qwen2.5-3B, if all feed-forward network (FFN) layers are replaced with MoE layers containing 16 experts, the resulting model’s non-embedding memory footprint will increase from 5.2GB to 73.2GB. This limitation is particularly pronounced for consumer-grade GPUs, which often have constrained memory capacities.
@@
Model quantization is a promising approach to reducing the memory footprint of LMMs while maintaining comparable performance. Most mainstream quantization methods FAHA (22); LTT+ (24); CCKDS (24); TSHDS (24) aim to compress the bit-width of a pre-trained, full-precision model. Although these methods have a low training cost, they suffer from significant performance degradation when the bit-width is reduced below 4-bit. Recent studies MWM+ (24); KVM+ (24); ZZS+ (24) have demonstrated promising scaling trends for ternary pre-training in Large Language Models (LLMs). At sufficiently large model sizes, ternary models can achieve accuracy comparable to full-precision models on downstream tasks while maintaining the same pre-training cost. Furthermore, they have much lower inference costs in terms of memory, latency, and energy consumption WZS+ (24). However, since these models have only been trained on billions of tokens, a substantial performance gap remains between open-sourced ternary models and full-precision dense models. As a result, directly training MoE models initialized from these under-trained models leads to weak performance on end tasks.
@@
In this work, we introduce MoTE, a scalable and memory-efficient architecture designed to train Mixture-of-Ternary Experts model from a pre-trained, full-precision dense checkpoint in multimodal tuning. Our approach addresses the inefficiency of multimodal MoE models in terms of memory footprint. Prior works LTY+ (24); LJH+ (25) primarily replace the FFN layer in dense checkpoints with an MoE layer, initializing the experts using the pre-trained FFN. However, we observed that in ternary training, replacing the FFN layer leads to significant performance degradation, as weight ternarization disrupts the pre-trained FFN. To mitigate this, we retain the FFN from the dense checkpoint as a shared expert activated for all inputs. During up-cycling, the layers inherited from the dense model remain frozen, while only the ternary MoE layers are trainable.
@@
We first conduct strict and controlled experiments to evaluate the proposed approach against full-precision up-cycling MoE-LLaVA LTY+ (24) across various model scales on a wide range of image understanding tasks. Our results show that ternary up-cycling exhibits surprising effectiveness as model size scales. As the size of the up-cycled dense checkpoint increases, the performance gap between MoTE and MoE-LLaVA narrows, eventually reaching comparable performance at scales larger than 1.5 billion parameters. Additionally, MoTE is compatible with post-training quantization techniques FAHA (22). Given the same expert memory footprint and combined with post-training quantization, MoTE outperforms full-precision MoE-LLaVA at both 1.5B and 3B model sizes. This advantage becomes even more pronounced as memory constraints tighten. Specifically, under an expert memory budget of 3.4GB, our approach achieves a 4.3% improvement in average accuracy on downstream task. These results demonstrate that given the same amount of total memory footprint and active parameter counts, training with a larger number of low-precision experts yields better performance than using fewer high-precision experts.
@@ 2 Related Work
LMMs demonstrate superior performance across various tasks as model size and training data scale increase. MoE models LLX+ (21); FZS (22); MSG+ (24); WCP+ (24) maintain a large pool of experts but activate only a subset for each token, enabling improved performance at the same FLOPs budget. KPL+ (23) introduced sparse up-cycling to reduce the training costs of MoE models by initializing them from dense checkpoints. LTY+ (24) explored the up-cycling of LMMs in the context of multimodal training, while SLZ+ (24) proposed a progressive knowledge transfer strategy to train small-scale multimodal MoEs from dense models. LJH+ (25) presented a scalable multimodal model that utilizes MoE with modality-specific encoders. While previous LTY+ (24); LWZ+ (24); LJH+ (25) primarily focused on full-precision experts for up-cycling, our work investigates up-cycling with ternary experts to develop memory-efficient multimodal MoE models.
@@
Quantization is a promising approach to reducing the memory footprint of LMMs while maintaining competitive performance, which can be categorized into two types based on the stage at which it is applied: post-training DLBZ (22); FAHA (22); LTT+ (24); CCKDS (24); TCS+ (24); TSHDS (24) and pre-training quantization WMD+ (23); MWM+ (24); WGL+ (25); PWW+ (23). Post-training quantization compresses high-precision pre-trained models after training. Due to its lower cost, it is widely adopted for mainstream large-scale models. GPTQ (FAHA, 22) and AWQ (LTT+, 24) reduce the bit-width to 4 bits while incurring minimal degradation. QuIP# (TCS+, 24) builds on QuIP (CCKDS, 24) by improving incoherence processing and applying vector quantization to incoherent weights. With additional fine-tuning, QuIP# achieves state-of-the-art performance in 2-bit models. However, when the bit-width is reduced below 4-bit, these methods all suffer from significant performance degradation compared to BF16 baselines. In contrast, pre-training quantization integrates quantization into the training process, requiring models to be trained from scratch, which results in better performance. Recent MWM+ (24) showed that ternary LLMs match the performance of full-precision counterpart starting from 3B parameter counts. FA (24) quantized a 1.6 trillion parameter Switch Transformer to sub 1-bit precision. (LJCC, 24) proposed to quantize the experts with a mixed precision recipe and introduced a novel data-driven techniques for optimizing bit allocation.
@@ 3.1 Architecture
We illustrate the architecture of MoTE in Figure 1. Previous studies KPL+ (23); LTY+ (24) expanded a dense model into an MoE model by directly replacing the FFN layer with an MoE layer, where each expert is initialized from the dense FFN to accelerate convergence. However, as shown in Table 6, we found that directly replacing the FFN with an MoE in ternary up-cycling leads to significant performance degradation. We hypothesize that this occurs because the FFN encodes a substantial amount of factual knowledge acquired during pre-training GSBL (21); DDH+ (22), and weight ternarization severely disrupts pre-trained information. To mitigate this issue, we retain the FFN module from the dense model as a shared expert, ensuring it is activated for every token. Specifically, the forward computation of the $l$-th layer of MoTE can be formulated as:
@@
[…]
@@
$\sigma$ is SiLU function. We apply absmean quantizer and per-token absmax quantizer for weight and activation quantization in expert’s linear layers following BitNet MWM+ (24). Specifically, the quantization can be formulated as:
@@
The weight $W\in\mathcal{R}^{m\times n}$ is quantized into ternary values, i.e., $\{-1,0,1\}$. The activations $x$ are per-token quantized into 8-bit integers, i.e., $[-128,127]$. The output of ternary linear layer $Y$ is $Q_{w}(W)Q_{a}(x)$. During inference, we use the kernel from BitBlas WMC+ (24) to save the memory footprint and accelerate the inference. Despite ternary values results in 1.58-bit, i.e., $\log 3/\log 2$, BitBlas still stores and processes ternary weight in INT2 format since current GPUs are still based on binary system.
@@ 3.2 Training recipe
Following MoE-LLaVA LTY+ (24), the training of MoTE consists of three stages. In Stage I, we train a two-layer MLP connector to align the visual encoder and LLM. As for Stage II, we fine-tune the LLM and connector using more complex vision-language instruction data. In Stage III, we expand the dense model from Stage II to an MoE model with ternary experts. The visual encoder is frozen through the training process. As presented in Figure 1, during up-cycling, only ternary MoE layers are trainable, and the shared expert and MSA layers are frozen.
We adopt quantization-aware training for MoTE. The weights and activations are quantized into ternary and INT8 values on-the-fly. Since many operations in the quantization are no-differentiable, we deploy straight-through estimator BLC (13) for gradient approximation. The gradients are directly by-passing through non-differentiable functions, i.e., $\frac{\partial\mathcal{L}}{\partial W}=\frac{\partial\mathcal{L}}{\partial Q(W)}$ and $\frac{\partial\mathcal{L}}{\partial X}=\frac{\partial\mathcal{L}}{\partial Q(X)}$. The gradients and optimizer states are retained as full-precision.
@@ 3.3 Training objectives
To ease the expert load imbalance problem in MoE, we adopt an auxiliary loss following Switch Transformers FZS (22). Given a batch of training tokens $\mathbf{X}$, the balancing loss can be formulated as:
@@ 1.5B Model Up-cycling
We select MoE-LLaVA LTY+ (24) as the baseline. It adopts a similar three-stage MoE training recipe and utilizes full-precision experts. Since MoE-LLaVA activates the top-2 experts, and our model includes a shared expert, we use top-1 gating in MoTE to ensure a fair comparison in terms of FLOPs. All MoE layers consist of four routed experts. We adopt SigLIP-L ZMKB (23) as the vision encoder and the instruct-version of Qwen2.5-series model YYZ+24a as the base LLM. The connector is a two-layer MLP with GELU activation. Table 1 presents the active and total parameter counts in the training of MoTE and MoE-LLaVA across different model sizes. The expert memory footprint includes contributions from both shared and routed experts.
@@
We adopt expert parallelism for efficient training of MoE models. The coefficient $\gamma$ for load balancing loss is set as 0.01. The value is recommended by FZS (22) to ensure auxiliary loss not to overwhelm the primary language modeling objective. All experiments are conducted on 16 NVIDIA A100 cards with 40GB memory. Due to the limited computation resources, we do not perform dynamic resolution processing for the images, since it leads to extremely long training sequence. The length of the total sequence is set as 2048 tokens, and the visual input includes 729 tokens. More hyper-parameters can be found in Appendix B.
@@
We train MoTE and MoE-LLaVA on the same dataset to ensure a fair comparison. The training dataset consists of a total of 5 million samples. For the first stage, we use the pre-training data of LLaVA 1.5 LLLL (24). For the second stage, we use the mixture of SViT ZWH (23), LVIS WMW+ (23), LRV LLL+ (23) and MIMIC-IT LZC+ (23). For the third stage, we use a subset of MAmmoTH-VL GZB+ (24), which includes 3.4 million instruction-response pairs, each associated with a single image as the visual input.
@@
We report the zero-shot performance of these models on a range of image understanding tasks using LMM-Eval toolkit ZLZ+ (24), including MMMU YNZ+ (24), MathVista LBX+ (24) (MathV), MMBench LDZ+ (24) (MMB), MMStar CLD+ (24) (MMS), MM-Vet YYL+ (23) (MMV), SeedBench-2-Plus LGC+ (24) (Seed2+), SeedBench LWW+ (23) (Seed), AI2D KSK+ (16), ChartQA MLT+ (22), InfoVQA MBT+ (22) and DocVQA MKJ (21).
@@ 4.2 Main results
We compared the performance of ternary up-cycling MoTE to MoE-LLaVA across different model sizes on various multimodal tasks. As shown in Table 2, MoTE underperformed full-precision up-cycling MoE-LLaVA when converting a 0.5B dense model to an MoE model. However, the performance gap between MoTE and MoE-LLaVA narrows as the parameter counts of the dense model increases. Similar phenomenons are also reported by the low-bit pre-training of LLMs WMD+ (23); MWM+ (24); KVM+ (24), which suggests promising trends of scaling model size for ternary MoEs.
@@ 1.5B Model Up-cycling
Model Training Tokens MMMU (val) MMB (en test) Seed (image) MMS (test) MMV (test) MathV (testmini) Avg.$\uparrow$
@@
MM1.5-1B ZGG+ (24) $>$200B 35.8 - 70.2 - 37.4 37.2 -
MM1.5-3B ZGG+ (24) $>$200B 37.1 - 72.4 - 41.0 44.4 -
MiniCPM-V2-3B YYZ+24b - 38.2 69.1 - 41.7 - 38.7 -
TinyLLaVA-3B ZHW+ (24) 4B 39.9 - - - 34.8 - -
Phi-3-Vision-4B AAA+ (24) $>$0.8T 40.4 73.9 71.8 47.9 45.4 44.5 54.0
Qwen2-VL-2B WBT+ (24) $>$1.4T 41.1 74.9 72.1 48.0 49.5 43.0 54.8
@@
MoE-LLaVA LTY+ (24) 4B 33.9 52.6 64.8 32.5 32.3 25.6 40.3
MolmoE-1B DCL+ (24)  1.5B 34.9 63.6 68.7 43.3 38.5 34.0 47.2
LLaVA-MoD-2B SLZ+ (24) 10B - 68.9 - - - - -
MM1-3B-MoE MGF+ (24) $>$400B 38.6 70.8 69.4 - 42.2 32.6 -
MM1-7B-MoE MGF+ (24) $>$400B 40.9 72.7 70.9 - 45.2 40.9 -
MM1.5-1B-MoE ZGG+ (24) $>$200B 41.2 - 71.4 - 39.8 42.9 -

### v2

@@ 1 Introduction
Large Multimodal Models (LMMs) Abdin et al. (2024); McKinzie et al. (2024); Zhang et al. (2024a); Wang et al. (2024c); Chen et al. (2024b); Bai et al. (2025) have achieved remarkable performance across a wide range of downstream tasks, including visual question answering and autonomous computer agents. However, as model size increases, the rising inference cost presents significant challenges for deploying LMMs efficiently. To address this, Mixture-of-Experts (MoE) Lepikhin et al. (2021); Fedus et al. (2022); DeepSeek-AI et al. (2024) introduces a mechanism that maintains a large pool of experts while activating only a subset for each input, thereby improving computational efficiency. Although MoE models significantly reduce FLOPs, they generally have a higher memory footprint, making deployment on edge devices challenging. For example, when training multimodal MoE up-cycled from Qwen2.5-3B, if all feed-forward network (FFN) layers are replaced with MoE layers containing 16 experts, the resulting model’s non-embedding memory footprint will increase from 5.2GB to 73.2GB. This limitation is particularly pronounced for consumer-grade GPUs, which often have constrained memory capacities.
@@
Model quantization is a promising approach to reducing the memory footprint of LMMs while maintaining comparable performance. Most mainstream quantization methods Frantar et al. (2022); Lin et al. (2024b); Chee et al. (2024); Tseng et al. (2024b) aim to compress the bit-width of a pre-trained, full-precision model. Although these methods have a low training cost, they suffer from significant performance degradation when the bit-width is reduced below 4-bit. Recent studies Ma et al. (2024); Kaushal et al. (2024); Zhu et al. (2024) have demonstrated promising scaling trends for ternary pre-training in Large Language Models (LLMs). At sufficiently large model sizes, ternary models can achieve accuracy comparable to full-precision models on downstream tasks while maintaining the same pre-training cost. Furthermore, they have much lower inference costs in terms of memory, latency, and energy consumption Wang et al. (2024a). However, since these models have only been trained on billions of tokens, a substantial performance gap remains between open-sourced ternary models and full-precision dense models. As a result, directly training MoE models initialized from these under-trained models leads to weak performance on end tasks.
@@
In this work, we introduce MoTE, a scalable and memory-efficient architecture designed to train Mixture-of-Ternary Experts model from a pre-trained, full-precision dense checkpoint in multimodal tuning. Our approach addresses the inefficiency of multimodal MoE models in terms of memory footprint. Prior works Lin et al. (2024a); Li et al. (2025) primarily replace the FFN layer in dense checkpoints with an MoE layer, initializing the experts using the pre-trained FFN. However, we observed that in ternary training, replacing the FFN layer leads to significant performance degradation, as weight ternarization disrupts the pre-trained FFN. To mitigate this, we retain the FFN from the dense checkpoint as a shared expert activated for all inputs. During up-cycling, the layers inherited from the dense model remain frozen, while only the ternary MoE layers are trainable.
@@
We first conduct strict and controlled experiments to evaluate the proposed approach against full-precision up-cycling MoE-LLaVA Lin et al. (2024a) across various model scales on a wide range of image understanding tasks. Our results show that ternary up-cycling exhibits surprising effectiveness as model size scales. As the size of the up-cycled dense checkpoint increases, the performance gap between MoTE and MoE-LLaVA narrows, eventually reaching comparable performance at scales larger than 1.5 billion parameters. Additionally, MoTE is compatible with post-training quantization techniques Frantar et al. (2022). Given the same expert memory footprint and combined with post-training quantization, MoTE outperforms full-precision MoE-LLaVA at both 1.5B and 3B model sizes. This advantage becomes even more pronounced as memory constraints tighten. Specifically, under an expert memory budget of 3.4GB, our approach achieves a 4.3% improvement in average accuracy on downstream task. These results demonstrate that given the same amount of total memory footprint and active parameter counts, training with a larger number of low-precision experts yields better performance than using fewer high-precision experts.
@@ 2 Related Work
LMMs demonstrate superior performance across various tasks as model size and training data scale increase. MoE models Lepikhin et al. (2021); Fedus et al. (2022); Muennighoff et al. (2024); Wu et al. (2024) maintain a large pool of experts but activate only a subset for each token, enabling improved performance at the same FLOPs budget. Komatsuzaki et al. (2023) introduced sparse up-cycling to reduce the training costs of MoE models by initializing them from dense checkpoints. Lin et al. (2024a) explored the up-cycling of LMMs in the context of multimodal training, while Shu et al. (2024) proposed a progressive knowledge transfer strategy to train small-scale multimodal MoEs from dense models. Li et al. (2025) presented a scalable multimodal model that utilizes MoE with modality-specific encoders. While previous Lin et al. (2024a); Li et al. (2024b, 2025) primarily focused on full-precision experts for up-cycling, our work investigates up-cycling with ternary experts to develop memory-efficient multimodal MoE models.
@@
Quantization is a promising approach to reducing the memory footprint of LMMs while maintaining competitive performance, which can be categorized into two types based on the stage at which it is applied: post-training Dettmers et al. (2022); Frantar et al. (2022); Lin et al. (2024b); Chee et al. (2024); Tseng et al. (2024a, b) and pre-training quantization Wang et al. (2023a); Ma et al. (2024); Wang et al. (2025); Peng et al. (2023). Post-training quantization compresses high-precision pre-trained models after training. Due to its lower cost, it is widely adopted for mainstream large-scale models. GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024b) reduce the bit-width to 4 bits while incurring minimal degradation. QuIP# (Tseng et al., 2024a) builds on QuIP (Chee et al., 2024) by improving incoherence processing and applying vector quantization to incoherent weights. With additional fine-tuning, QuIP# achieves state-of-the-art performance in 2-bit models. However, when the bit-width is reduced below 4-bit, these methods all suffer from significant performance degradation compared to BF16 baselines. In contrast, pre-training quantization integrates quantization into the training process, requiring models to be trained from scratch, which results in better performance. Recent Ma et al. (2024) showed that ternary LLMs match the performance of full-precision counterpart starting from 3B parameter counts. Frantar and Alistarh (2024) quantized a 1.6 trillion parameter Switch Transformer to sub 1-bit precision. (Li et al., 2024c) proposed to quantize the experts with a mixed precision recipe and introduced a novel data-driven techniques for optimizing bit allocation.
@@ 3.1 Architecture
We illustrate the architecture of MoTE in Figure 1. Previous studies Komatsuzaki et al. (2023); Lin et al. (2024a) expanded a dense model into an MoE model by directly replacing the FFN layer with an MoE layer, where each expert is initialized from the dense FFN to accelerate convergence. However, as shown in Table 6, we found that directly replacing the FFN with an MoE in ternary up-cycling leads to significant performance degradation. We hypothesize that this occurs because the FFN encodes a substantial amount of factual knowledge acquired during pre-training Geva et al. (2021); Dai et al. (2022), and weight ternarization severely disrupts pre-trained information. To mitigate this issue, we retain the FFN module from the dense model as a shared expert, ensuring it is activated for every token. Specifically, the forward computation of the $l$-th layer of MoTE can be formulated as:
@@
where MSA and LN stands for multi-head self-attention and layer normalization, respectively. As illustrated in Figure 1, we initialize the FFN, MSA and MoE layers from the dense model. We implement the MoE mechanism following the GShard Lepikhin et al. (2021), with each expert modeled as a Gated Linear Unit (GLU) Shazeer (2020). An MoE layer which consists of $E$ ternary experts $\text{FFN}^{T}_{1}$$\text{FFN}^{T}_{E}$ satisfies that:
@@
$\sigma$ is SiLU function. We apply absmean quantizer and per-token absmax quantizer for weight and activation quantization in expert’s linear layers following BitNet Ma et al. (2024). Specifically, the quantization can be formulated as:
@@
The weight $W\in\mathcal{R}^{m\times n}$ is quantized into ternary values, i.e., $\{-1,0,1\}$. The activations $x$ are per-token quantized into 8-bit integers, i.e., [128,127][-128,127]. The output of ternary linear layer $Y$ is $Q_{w}(W)Q_{a}(x)$. During inference, we use the kernel from BitBlas Wang et al. (2024b) to save the memory footprint and accelerate the inference. Despite ternary values results in 1.58-bit, i.e., $\log 3/\log 2$, BitBlas still stores and processes ternary weight in INT2 format since current GPUs are still based on binary system.
@@ 3.2 Training recipe
Following MoE-LLaVA Lin et al. (2024a), the training of MoTE consists of three stages. In Stage I, we train a two-layer MLP connector to align the visual encoder and LLM. As for Stage II, we fine-tune the LLM and connector using more complex vision-language instruction data. In Stage III, we expand the dense model from Stage II to an MoE model with ternary experts. The visual encoder is frozen through the training process. As presented in Figure 1, during up-cycling, only ternary MoE layers are trainable, and the shared expert and MSA layers are frozen.
We adopt quantization-aware training for MoTE. The weights and activations are quantized into ternary and INT8 values on-the-fly. Since many operations in the quantization are no-differentiable, we deploy straight-through estimator Bengio et al. (2013) for gradient approximation. The gradients are directly by-passing through non-differentiable functions, i.e., $\frac{\partial\mathcal{L}}{\partial W}=\frac{\partial\mathcal{L}}{\partial Q(W)}$ and $\frac{\partial\mathcal{L}}{\partial X}=\frac{\partial\mathcal{L}}{\partial Q(X)}$. The gradients and optimizer states are retained as full-precision.
@@ 3.3 Training objectives
To ease the expert load imbalance problem in MoE, we adopt an auxiliary loss following Switch Transformers Fedus et al. (2022). Given a batch of training tokens $\mathbf{X}$, the balancing loss can be formulated as:
@@ 1.5B Model Up-cycling
We select MoE-LLaVA Lin et al. (2024a) as the baseline. It adopts a similar three-stage MoE training recipe and utilizes full-precision experts. Since MoE-LLaVA activates the top-2 experts, and our model includes a shared expert, we use top-1 gating in MoTE to ensure a fair comparison in terms of FLOPs. All MoE layers consist of four routed experts. We adopt SigLIP-L Zhai et al. (2023) as the vision encoder and the instruct-version of Qwen2.5-series model Yang et al. (2024) as the base LLM. The connector is a two-layer MLP with GELU activation. Table 1 presents the active and total parameter counts in the training of MoTE and MoE-LLaVA across different model sizes. The expert memory footprint includes contributions from both shared and routed experts.
@@
We adopt expert parallelism for efficient training of MoE models. The coefficient $\gamma$ for load balancing loss is set as 0.01. The value is recommended by Fedus et al. (2022) to ensure auxiliary loss not to overwhelm the primary language modeling objective. Due to the limited computation resources, we do not perform dynamic resolution processing for the images, since it leads to extremely long training sequence. The length of the total sequence is set as 2048 tokens, and the visual input includes 729 tokens. More hyper-parameters can be found in Appendix B.
@@
We train MoTE and MoE-LLaVA on the same dataset to ensure a fair comparison. The training dataset consists of a total of 5 million samples. For the first stage, we use the pre-training data of LLaVA 1.5 Liu et al. (2024a). For the second stage, we use the mixture of SViT Zhao et al. (2023), LVIS Wang et al. (2023b), LRV Liu et al. (2023) and MIMIC-IT Li et al. (2023a). For the third stage, we use a subset of MAmmoTH-VL Guo et al. (2024), which includes 3.4 million instruction-response pairs, each associated with a single image as the visual input.
@@
We report the zero-shot performance of these models on a range of image understanding tasks using LMM-Eval toolkit Zhang et al. (2024b), including MMMU Yue et al. (2024), MathVista Lu et al. (2024) (MathV), MMBench Liu et al. (2024b) (MMB), MMStar Chen et al. (2024a) (MMS), MM-Vet Yu et al. (2023) (MMV), SeedBench-2-Plus Li et al. (2024a) (Seed2+), SeedBench Li et al. (2023b) (Seed), AI2D Kembhavi et al. (2016), ChartQA Masry et al. (2022), InfoVQA Mathew et al. (2022) and DocVQA Mathew et al. (2021).
@@ 4.2 Main results
We compared the performance of ternary up-cycling MoTE to MoE-LLaVA across different model sizes on various multimodal tasks. As shown in Table 2, MoTE underperformed full-precision up-cycling MoE-LLaVA when converting a 0.5B dense model to an MoE model. However, the performance gap between MoTE and MoE-LLaVA narrows as the parameter counts of the dense model increases. Similar phenomenons are also reported by the low-bit pre-training of LLMs Wang et al. (2023a); Ma et al. (2024); Kaushal et al. (2024), which suggests promising trends of scaling model size for ternary MoEs.
@@ 1.5B Model Up-cycling
[…]
@@
MM1.5-1B Zhang et al. (2024a) >>200B 35.8 - 70.2 - 37.4 37.2 -
MM1.5-3B Zhang et al. (2024a) >>200B 37.1 - 72.4 - 41.0 44.4 -
MiniCPM-V2-3B Yao et al. (2024) - 38.2 69.1 - 41.7 - 38.7 -
TinyLLaVA-3B Zhou et al. (2024) 4B 39.9 - - - 34.8 - -
Phi-3-Vision-4B Abdin et al. (2024) >>0.8T 40.4 73.9 71.8 47.9 45.4 44.5 54.0
Qwen2-VL-2B Wang et al. (2024c) >>1.4T 41.1 74.9 72.1 48.0 49.5 43.0 54.8
@@
MoE-LLaVA Lin et al. (2024a) 4B 33.9 52.6 64.8 32.5 32.3 25.6 40.3
MolmoE-1B Deitke et al. (2024)  1.5B 34.9 63.6 68.7 43.3 38.5 34.0 47.2
LLaVA-MoD-2B Shu et al. (2024) 10B - 68.9 - - - - -
MM1-3B-MoE McKinzie et al. (2024) >>400B 38.6 70.8 69.4 - 42.2 32.6 -
MM1-7B-MoE McKinzie et al. (2024) >>400B 40.9 72.7 70.9 - 45.2 40.9 -
MM1.5-1B-MoE Zhang et al. (2024a) >>200B 41.2 - 71.4 - 39.8 42.9 -